OpenStudy (anonymous):

what is 2yards to 45 feet as a fraction

OpenStudy (anonymous):

Help me !

OpenStudy (jack1):

how many feet in a yard? (damn ur imperial system, switch to metric already, who uses fathoms anyway) ;D

OpenStudy (jack1):

srsly @kyanna_Hester how many feet in a yard, i dont know, can u tell me and ill help u

OpenStudy (anonymous):

3 feet is in a yard

OpenStudy (jack1):

ok so 2 yards : 45 ft so 6 ft : 45 ft so 6/45 so 2/15

OpenStudy (anonymous):

30 mi / h = ____ ft / min ?

OpenStudy (jack1):

again: non-metric how many miles in a gallon or whatever you want to convert it to...? feet was it?

OpenStudy (anonymous):

yea

OpenStudy (jack1):

k how many feet in a mile?

OpenStudy (anonymous):

5,280

OpenStudy (jack1):

\[\frac{ 30miles }{1 hours }\times \frac{ 5280 ft }{ 1 mile }\times \frac{ 1hour }{ 60 mins } =answer\]

OpenStudy (jack1):

so times the top line across =...? then times the bottom line across =...?

OpenStudy (anonymous):

whats is the answer?

OpenStudy (jack1):

u tell me

OpenStudy (anonymous):

idk ?

OpenStudy (jack1):

so times the top line across =...30 x 5280 x 1 =?

OpenStudy (anonymous):

Express each ratio as a fraction in simplest form. 34 laptops out of 116 computers

OpenStudy (jack1):

wait... what was the answer to the previous q...?

OpenStudy (anonymous):

my teacher said you were wrong?! so next question ig

OpenStudy (jack1):

then ur teacher is wrong top line = 30 x 5280 x 1= 158,400 ft bottom line = 1 x 1 x 60 = 60 mins answer = 158400 ft / 60mins = 2640 ft/min

OpenStudy (anonymous):

lol i showed him your answer and you made him look stupid ! haha ok whats..... Express each ratio as a fraction in simplest form. 34 laptops out of 116 computers

OpenStudy (jack1):

what are the factors of 34? 1x34 2x17 now what are the factors of 116 1x116 2x 58 4x 29

OpenStudy (jack1):

so to match the common factors: both are divisible by 2 only 34/2 = 17 116/2 =58 ans = 17/58

OpenStudy (anonymous):

Convert each rate using dimensional analysis. Round to the nearest tenth, if necessary. 55 mi / h = ____ ft / s

OpenStudy (anonymous):

my teacher said why does it take you so long?!